# Negative non-Gaussianity as a salvager for PBHs with PTAs in bounce

### v1

@@ VII Comparative analysis of previous studies
This section aims to address previous analyses conducted in line with studying the overproduction of PBHs and compares them through their efficiency and limitations in contrast to the approach adopted for this work and its results. Very recently, since the definitive announcement of an SGWB signal via the PTA collaborations, several groups have highlighted a dramatic issue concerning the abundant production of PBHs. This issue directly relates to having the SIGWs as a leading interpretation of the PTA signal. Upon its confrontation with the observed signal, it suggests a highly amplified scalar power spectrum, which implies that the PBHs can get overproduced. Below we discuss some of the major advances specific to this problem and supply the table [^1] Note that the table serves to display a certain parameters of few chosen models and no attempt has been made to present a comprehensive list of all the models studied so far.1 summarizing the same.
@@
The emergence of QCD effects in the PTA signal stems from the sudden decrease in the energy and entropy degrees of freedom around the temperature, $T\gtrsim 100{\rm MeV}$, that aligns perfectly well within the regime of frequency probed by the PTA and thus the corresponding decrease in the equation of state can noticeably alter the low-frequency tail of the signal [489]. Using such conditions for Galileon inflation, [20] shows the softening of the IR tail of the GW spectrum till $f\simeq 5\times 10^{-9}{\rm Hz}$, that is, before we reach the peak of the spectrum. In the present study however, QCD effects have not been taken into account for formation of PBHs.
In the last row, we consider the case where quantum loop effects are considered while estimating the results on PBH production from a model of interest. In the light of the no-go theorem [233, 234, 235], all the studies mentioned, except for the curvaton, come as examples of how to successfully evade the stringent bound, $M_{\rm PBH}\sim{\cal O}(10^{2}{\rm gms})$, as a consequence of this theorem.
The present work applies the same spirit of using the PNGs to our advantage but goes a step further by taking into account enormous values for the same, $f_{\rm NL}\ll{\cal O}(-10)$, in comparison to the previous studies. This work also exploits the compaction function approach rather than the standard Press-Schechter formalism or the method from peaks theory to study the PBH formation process. And at last, the underlying set-up of this paper goes beyond the choice of working just within the inflationary picture, that numerous other studies above prefer to address this issue, but also combines the features and consequences of having a non-singular bouncing scenario before the usual slow-roll commences. To conduct the analysis for the overproduction problem, with section VI.3 devoted towards its end results, we utilize the theoretical set-up that is built upon the works [233, 234, 235, 240] and by integrating the aforementioned improvements we aim to achieve a stronger and broader understanding on curing the overproduction along with shedding some light in an upcoming section on the acceptable amounts of such type of NGs.
@@ VIII Confrontation with experiments
After conducting an elaborate discussion including various models and their relationship with multiple parameters encountered during PBH formation, we briefly discuss in this section their results when confronted with different gravitational wave experiments and, if possible, microlensing experiments for searching signatures of dark matter. The summary of these results can be observed through table 2.
@@ IX Establishment of strong upper bound on $f_{\rm NL}$
The need for amplified strength of inflationary fluctuations for PBH formation does not necessarily warrant the validation of perturbativity constraints. From the figure (6), we observe a monotonic trend for the peak amplitude $A$ with increasing magnitude of the NG parameter $f_{\rm NL}$. Our previous results have shown how helpful large negative NGs are in dealing with PBH overproduction. Presently, in both figures 6 and 6, we can see that the amplitude $A$ required to achieve $f_{\rm PBH}\sim 1$, increases with negative $f_{\rm NL}$ as well as with higher values of the compaction threshold ${\cal C}_{\rm th}$. However, there comes a moment where the required amplitude crosses over $A\sim{\cal O}(0.1)$ for both $f_{\rm NL}=(-39.95,-35/8)$ and $c_{s}=1$ and over $A\sim{\cal O}(0.01)$ for the same set of $f_{\rm NL}$ but now with $c_{s}=0.88$. Following our repeated emphasis over choosing the canonical case where $c_{s}=1$, we notice from figure 6 that going beyond $f_{\rm NL}\sim-60$ will surely keep pushing the amplitude close of eventually hit $A\sim{\cal O}(1)$ and putting the analysis at risk of breaking perturbativity. In a similar spirit, with $c_{s}=0.88$ in figure 6, one observes the condition $A>0.01$ for $f_{\rm NL}\lesssim-60$. Furthermore, since we aim for the interval, $f_{\rm PBH}\in(10^{-3},1)$, that quantifies as having a sizeable fraction, the maximum amplitude difference tolerable for a particular wavenumber comes to be, $\log_{10}(\Delta A)=-0.2$, which is an estimate that we can expect even when we consider $f_{\rm NL}\sim{\cal O}(-60)$.
@@
For the space of allowed compaction thresholds, the values that we have considered here for ${\cal C}_{\rm th}$ lie in between $(0.5,0.65)$, with ${\cal C}_{\rm th}=0.5$ demanding the amplitude to come just close to $A\sim 0.1$. And so, any value of this threshold lower than ${\cal C}_{\rm th}\sim 0.5$ will accordingly require a low enough amplitude even if we consider $f_{\rm NL}\sim-60$. Thus, we conclude that for $0.50.4<{\cal C}_{\rm th}<$ 0.5, where the lower limit arises from our choice of working with the linear regime result, ${\cal C}_{\rm th}\in[2/5,2/6]$ [461], the strength of the scalar power spectrum is still sufficient to incorporate further large and negative NG, $f_{\rm NL}<-60$. The significance of $f_{\rm NL}=-60$ also extends to the fact that in the figure (6) we choose to show the analysis for the same set of compaction thresholds that we previously incorporate in figures (3-5) to highlight the removal of PBH overproduction with the choice of ${\cal C}_{\rm th}\in(0.5,0.65)$ giving us the best agreement with the data.
@@
Considering the pre-inflationary dynamics of the perturbations, studies have pointed out that, in principle, one great advantage of such conditions is the generation of a large amount of NG that reaches upto $f_{\rm NL}\sim{\cal O}(100)$ [302, 480]. In such cases, we can, however, comment about large NGs as our use of an EFT framework encompasses both the scenarios before the beginning of inflation and hereafter, thus allowing for a generalized study. Hence, concluding from the observations, we provide a stringent upper bound on the possible magnitude of NG, $f_{\rm NL}\sim-60$, that does not endanger the validity of our analysis within our EFT setup.
@@ X Conclusion
In this section, we summarize our crucial findings regarding the cure for PBH overproduction and its connection with exhibiting large and negative primordial non-Gaussianities. Our underlying set-up takes advantage of the features coming from pre-inflationary physics, with the introduction of the contracting and bouncing phases, into the generation of extremely large non-Gaussianities in the curvature perturbations. The benefit of bearing such large NGs is reflected in controlling the abundance of PBHs, which are formed during inflation, and they determine whether their abundance gets either suppressed or enhanced. We began with a brief introduction on the EFT of bounce framework built on the works of [233, 234, 235, 240], and described how the scalar power spectrum after contributions from each phase gets constructed including the one-loop quantum corrections. In order to provide a lucid procedure of handling the quantum-loop corrections, we summarize its main components of regularization, renormalization, and resummation which altogether assist in removing any quadratic (or power-law) UV and smoothen the logarithmic IR divergences resulting from each phase to the tree level result. At the end of the procedures, we obtained the crucial one-loop renormalized and DRG resummed scalar power spectrum through which we later discuss the production of PBHs.
Keeping in line with the latest surge in the theoretical advancements of PBH formation, we preferred to employ the compaction function criterion for the same and elaborate on its formalism which provides a means to calculate the PBH mass fraction. We further integrate the inclusion of quadratic form of NGs within the compaction function approach, to conduct a qualitative study of the impact NGs have on PBH formation. We follow this with a discussion on how we go about calculating the necessary abundance and the effect of having PNGs in such calculations. After this, we take a brief interlude to discuss the energy density spectrum of the scalar induced gravity waves in the radiation dominated era where $w=1/3$. Throughout this work, we do not entertain any free equation of state and proceed with the fact that the scales contributing most to the PBH collapse do so after their re-entry into the radiation era, for a general equation of state analysis within the EFT of bounce set-up we refer to [303].
@@
    We selected two distinct benchmark values for investigating PNGs namely, $f_{\rm NL}=(-39.95,-35/8)$, which result from independent analysis focusing on either an ekpyrotic contraction [302] or an expanding period of matter bounce [301], respectively. Within this analysis, we added another variable - effective sound speed component and considered its two separate values, $c_{s}=(0.88,1)$ based on a previous work [240].
@@
    To highlight our best result, we established that larger negative values like $f_{\rm NL}=-39.95$, coupled with the canonical case $c_{s}=1$, proved utmost beneficial in increasing the abundance of PBHs and simultaneously establishing an excellent agreement of the requisite peak power spectrum amplitude within $1\sigma$ of having a SIGW explanation of the recent PTA signal.
@@
    Utilizing a lower value of $c_{s}=0.88$ lead to an order of magnitude suppression in the amplitude, pushing the agreement to outside the $2\sigma$ regime when $f_{\rm NL}=-39.95$ and even outside by $3\sigma$ for the choice of $f_{\rm NL}=-38/5$.
@@
    After this, we conducted a comparative study highlighting the latest attempts towards addressing overproduction through various models that either include just the PNGs to their advantage or incorporate other new parameters. We followed this by confronting these models’ results with a few observational experiments for detecting gravitational wave signals and candidates for dark matter via microlensing.
@@
    Later, we scrutinized the acceptance of negatively large NGs in our EFT model and concluded this discussion by extracting an upper bound of $f_{\rm NL}\sim-60$ as the most allowable negative value for the case of large compaction thresholds, ${\cal C}_{\rm th}\in(0.5,0.65)$, utilizing which we earlier mitigated strongly the overproduction of PBHs.
@@
    When going more negative beyond $f_{\rm NL}\sim-60$, we found the power spectrum amplitude to cross over the condition $A\sim{\cal O}(0.1)$ for $f_{\rm NL}=-39.95$, and also over $A\sim{\cal O}(0.01)$ for $f_{\rm NL}=-35/8$, which signalled a breach in the underlying perturbativity arguments.
@@ Acknowledgement
Appendix A Comoving curvature perturbation mode solutions

### v3

@@ VII Comparative analysis of previous studies
This section aims to address previous analyses conducted in line with studying the overproduction of PBHs and compares them through their efficiency and limitations in contrast to the approach adopted for this work and its results. Quite recently, since the definitive announcement of an SGWB signal via the PTA collaborations, several groups have highlighted a dramatic issue concerning the abundant production of PBHs. This issue directly relates to the SIGWs being a leading interpretation of the PTA signal. Upon confrontation with the observed signal, it suggests a highly amplified scalar power spectrum, which implies that the PBHs can be overproduced. Below we discuss some of the major advances specific to this problem and supply the table [^1] Note that the table serves to display a certain parameters of few chosen models and no attempt has been made to present a comprehensive list of all the models studied so far.1 summarizing the same.
@@
The emergence of QCD effects in the PTA signal stems from the sudden decrease in the energy and entropy degrees of freedom around the temperature, $T\gtrsim 100{\rm MeV}$, which aligns perfectly within the frequency regime probed by the PTA and thus the corresponding decrease in the equation of state can noticeably alter the low-frequency tail of the signal [505]. Using such conditions for Galileon inflation, [20] shows the softening of the IR tail of the GW spectrum to $f\simeq 5\times 10^{-9}{\rm Hz}$, that is, before we reach the peak of the spectrum. However, in the present study, QCD effects have not been taken into account for the formation of PBHs.
In the last row, we consider the case where quantum loop effects are considered while estimating the results on PBH production from a model of interest. In light of the no-go theorem [244, 245, 246], all the studies mentioned, except for the curvaton, come as examples of how to successfully evade the stringent bound, $M_{\rm PBH}\sim{\cal O}(10^{2}{\rm gms})$, as a consequence of this theorem.
The present work applies the same spirit of using the PNGs to our advantage but goes a step further by taking into account enormous values for the same, $f_{\rm NL}\ll{\cal O}(-10)$, compared to previous studies. This work also exploits the compaction function approach rather than the standard Press-Schechter formalism or the method from the peaks theory to study the PBH formation process. And finally, the underlying setup of this paper goes beyond the choice of working just within the inflationary picture that numerous other studies above prefer to address this issue, but also combines the features and consequences of having a non-singular bouncing scenario before the usual slow-roll commences. To conduct the analysis for the overproduction problem, with section VI.3 devoted towards its end results, we utilize the theoretical set-up that is built upon the works [244, 245, 246, 253] and by integrating the aforementioned improvements we aim to achieve a stronger and broader understanding of curing the overproduction along with shedding some light in an upcoming section on the acceptable amounts of such type of NG.
@@ VIII Confrontation with experiments
After conducting an elaborate discussion including various models and their relationship with multiple parameters encountered during PBH formation, we briefly discuss in this section their results when confronted with different gravitational wave experiments and, if possible, microlensing experiments for searching signatures of dark matter. The summary of these results can be seen in Table 2.
@@ IX Establishment of strong upper bound on $f_{\rm NL}$
The need for amplified strength of inflationary fluctuations for PBH formation does not necessarily warrant the validation of perturbativity constraints. From the figure (6), we observe a monotonic trend for the peak amplitude $A$ with increasing magnitude of the NG parameter $f_{\rm NL}$. Our previous results have shown how helpful large negative NGs are in dealing with PBH overproduction. Presently, in both figures 6 and 6, we can see that the amplitude $A$ required to achieve $f_{\rm PBH}\sim 1$, increases with negative $f_{\rm NL}$ as well as with higher values of the compaction threshold ${\cal C}_{\rm th}$. However, there comes a moment where the required amplitude crosses over $A\sim{\cal O}(0.1)$ for both $f_{\rm NL}=(-39.95,-35/8)$ and $c_{s}=1$ and over $A\sim{\cal O}(0.01)$ for the same set of $f_{\rm NL}$ but now with $c_{s}=0.88$. Following our repeated emphasis over choosing the canonical case where $c_{s}=1$, we notice from figure 6 that going beyond $f_{\rm NL}\sim-60$ will surely keep pushing the amplitude close to eventually hitting $A\sim{\cal O}(1)$ and putting the analysis at risk of breaking perturbativity. In a similar spirit, with $c_{s}=0.88$ in figure 6, one observes the condition $A>0.01$ for $f_{\rm NL}\lesssim-60$. Furthermore, since we aim for the interval, $f_{\rm PBH}\in(10^{-3},1)$, which quantifies as having a sizeable fraction, the maximum amplitude difference tolerable for a particular wavenumber comes to, $\log_{10}(\Delta A)=-0.2$, which is an estimate that we can expect even when we consider $f_{\rm NL}\sim{\cal O}(-60)$.
@@
For the space of allowed compaction thresholds, the values we have considered here for ${\cal C}_{\rm th}$ lie between $(0.5,0.65)$, with ${\cal C}_{\rm th}=0.5$ demanding that the amplitude come just close to $A\sim 0.1$. So, any value of this threshold lower than ${\cal C}_{\rm th}\sim 0.5$ will accordingly require a low enough amplitude even if we consider $f_{\rm NL}\sim-60$. Thus, we conclude that for $0.50.4<{\cal C}_{\rm th}<$ 0.5, where the lower limit arises from our choice to work with the linear regime result, ${\cal C}_{\rm th}\in[2/5,2/6]$ [477], the strength of the scalar power spectrum is still sufficient to incorporate further large and negative NG, $f_{\rm NL}<-60$. The significance of $f_{\rm NL}=-60$ also extends to the fact that in figure (6) we choose to show the analysis for the same set of compaction thresholds that we previously incorporate in figures (3-5) to highlight the removal of PBH overproduction with the choice of ${\cal C}_{\rm th}\in(0.5,0.65)$, which gives us the best agreement with the data.
Some models also predict extremely large but positive NGs closer to the wavenumber regime where the power spectrum gets enhanced due to modifications in the inflationary potential, with one such study addressing this feature present in [509]. The model there supposes modifying the $\alpha-$attractor potential via a tiny bump to enhance the primordial power spectrum [510], which promotes the formation of PBHs. In [509], they show that at small scales, close to the wavenumbers that lead to PBHs, the value of the equilateral NG configuration reaches $f^{\rm equil}_{\rm NL}\sim{\cal O}(10^{3})$. In contrast to this feature, our study focuses on the implications of local type or squeezed NGs that are proven, as stated at the beginning of this section, to dominate models that have an ekpyrotic or matter-type contraction or bouncing phase. After choosing a set of $f_{\rm NL}$ values obtained within models with such pre-inflationary features, we conduct our study of curing PBH overproduction and finally propose the aforementioned upper bound on negatively large NGs.
@@
Considering the pre-inflationary dynamics of the perturbations, studies have pointed out that, in principle, one great advantage of such conditions is the generation of a large amount of NG that reaches $f_{\rm NL}\sim{\cal O}(-100)$ [318, 496]. In such cases, we can, however, comment about large NGs as our use of an EFT framework encompasses both the scenarios before the beginning of inflation and hereafter, thus allowing for a generalized study. Hence, upon conclusion from the observations, we provide a strict upper bound on the possible magnitude of NG, $f_{\rm NL}\sim-60$, which does not endanger the validity of our analysis within our EFT setup.
@@ X Conclusion
In this section, we summarize our crucial findings regarding the cure for PBH overproduction and its connection with exhibiting large and negative primordial non-Gaussianities. Our underlying set-up takes advantage of the features coming from pre-inflationary physics, with the introduction of the contracting and bouncing phases, into the generation of extremely large non-Gaussianities in the curvature perturbations. The benefit of bearing such large NGs is reflected in controlling the abundance of PBHs, which are formed during inflation, and they determine whether their abundance gets either suppressed or enhanced. We began with a brief introduction on the EFT of bounce framework built on the works of [244, 245, 246, 253], and described how the scalar power spectrum after contributions from each phase is constructed including the one-loop quantum corrections. In order to provide a lucid procedure of handling the quantum-loop corrections, we summarize its main components of regularization, renormalization, and resummation which altogether assist in removing any quadratic (or power-law) UV and smoothen the logarithmic IR divergences resulting from each phase to the tree-level result. At the end of the procedures, we obtained the crucial one-loop renormalized and DRG resummed scalar power spectrum through which we later discuss the production of PBHs.
Keeping in line with the latest surge in the theoretical advancements of PBH formation, we preferred to employ the compaction function criterion for the same and elaborate on its formalism, which provides a means to calculate the PBH mass fraction. We further integrate the inclusion of the quadratic form of NGs within the compaction function approach, to conduct a qualitative study of the impact NGs have on PBH formation. We follow this with a discussion on how we go about calculating the necessary abundance and the effect of having PNGs in such calculations. After this, we take a brief interlude to discuss the energy density spectrum of scalar-induced gravity waves in the radiation-dominated era where $w=1/3$. Throughout this work, we do not entertain any free equation of state and proceed with the fact that the scales contributing most to the PBH collapse do so after their re-entry into the radiation era, for a general equation of state analysis within the EFT of bounce set-up we refer to [319].
@@
    We selected two distinct benchmark values for investigating PNGs namely, $f_{\rm NL}=(-39.95,-35/8)$, which result from independent analysis focusing on either an ekpyrotic contraction [318] or an expanding period of matter bounce [317], respectively. Within this analysis, we added another variable, the effective sound speed component, and considered its two separate values, $c_{s}=(0.88,1)$ based on a previous work [253].
@@
    To highlight our best result, we established that larger negative values such as $f_{\rm NL}=-39.95$, coupled with the canonical case $c_{s}=1$, proved utmost beneficial in increasing the abundance of PBHs and simultaneously establishing an excellent agreement of the requisite peak power spectrum amplitude within $1\sigma$ of having a SIGW explanation of the recent PTA signal.
@@
    Utilizing a lower value of $c_{s}=0.88$ leads to an order-of-magnitude suppression in the amplitude, pushing the agreement outside the $2\sigma$ regime when $f_{\rm NL}=-39.95$ and even outside by $3\sigma$ for the choice of $f_{\rm NL}=-38/5$.
@@
    After this, we conducted a comparative study highlighting the latest attempts towards addressing overproduction through various models that include only the PNGs to their advantage or incorporate other new parameters. We followed this by confronting these models’ results with a few observational experiments for detecting gravitational wave signals and candidates for dark matter via microlensing.
@@
    Later, we scrutinized the acceptance of negatively large NGs in our EFT model and concluded this discussion by extracting an upper bound of $f_{\rm NL}\sim-60$ as the most allowable negative value for the case of large compaction thresholds, ${\cal C}_{\rm th}\in(0.5,0.65)$, utilizing which we previously strongly mitigated the overproduction of PBHs.
@@
    When going more negative beyond $f_{\rm NL}\sim-60$, we found the power spectrum amplitude to cross over the condition $A\sim{\cal O}(0.1)$ for $f_{\rm NL}=-39.95$, and also over $A\sim{\cal O}(0.01)$ for $f_{\rm NL}=-35/8$, signalling a breach in the underlying perturbativity arguments.
@@ Acknowledgement
Appendix A Evolution of slow-roll parameters in the EFT of bounce framework
In this section, we discuss the evolution of the two slow-roll parameters, $\epsilon(N)$ and $\eta(N)$, depending on the nature of the pre-inflationary setup via its corresponding scale factor, $a(t)$, and how such parameters then evolve throughout the complete EFT of bounce framework studied in this work. We use the two scale-factor definitions as stated previously in Eqs. (2-3).
In order to derive the behaviour of these slow-roll parameters, we would use the help of the simple relation, $dN=Hdt={\cal H}d\tau$, for the e-folding elapsed in a conformal time interval, $d\tau$, and where, ${\cal H}=aH$, is the conformal Hubble parameter. Let us first analyze the slow-roll parameters during the pre-inflationary phases, after which we will then enter the SR/USR/SR setup. The final result for the evolution is shown in the figures (7-8) for the complete setup.
  1. (a)
    Phase-I  (Contraction): This phase can be an ekpyrotic or matter-type contraction phase. Using the definition in Eqn. (2) we can write the relation, ${\cal H}=a^{\prime}/a=1/(\tau(\epsilon-1))$. Once we know this, we can now proceed with the following relation for the conformal time as:
[…]
    for $N<N_{c}$ where $N=N_{c}$ signals the end of the contraction phase, and $N_{0}$ and $\tau_{0}$ are reference variables for the e-folding and conformal time such that $a(\tau_{0})=a_{0}$ in Eqn. (2). The above result allows us to obtain the conformal Hubble parameter as a function of the e-folds $N$ during contraction. Now we are ready to write the relation for $\epsilon(N)$ using its following definition:
    $\epsilon(N)$ $=$ $1-\frac{{\cal H}^{\prime}}{{\cal H}^{2}}=1+\frac{\tau^{2}(\epsilon-1)^{2}}{\tau^{2}(\epsilon-1)}=\epsilon,$ (81)
    where $\epsilon$ in the RHS takes the value $\epsilon=3/2$ during matter contraction or $\epsilon>3$ during the ekpyrotic contraction phase. In a similar fashion, we can derive the relation with e-folding for the second slow-roll parameter $\eta(N)$. We use its corresponding definition to arrive at the following relation:
    $\eta=\frac{\epsilon^{\prime}}{\epsilon{\cal H}}=0,$ (82)
    since $\epsilon$ does not change during this phase.
    Figure 7: The first slow-roll parameter $\epsilon(N)$ as a function of the e-folds $N$ in 7 with the matter contraction (MCP) and matter bounce (MBP) phases and in 7 with the ekpyrotic contraction (ECP) and ekpkyrotic bounce (EBP) phases followed by the SRI, USR, and SRII phases.
    Figure 8: The second slow-roll parameter $\eta(N)$ as a function of the e-folds $N$ in 8 with the matter contraction (MCP) and matter bounce (MBP) phases and in 8 with the ekpyrotic contraction (ECP) and ekpkyrotic bounce (EBP) phases followed by the SRI, USR, and SRII phases.
  2. (b)
    Phase-II  (Bounce): This phase can also be an ekpyrotic or matter-type bouncing phase. This time after utilising the definition of the scale factor in Eqn. (3) we get the following relation for the conformal time as:
[…]
    $\implies N-N_{0}$ $=$ $\frac{\ln{[\tau^{2}+\tau_{0}^{2}]}}{2(\epsilon-1)}\Biggr{|}_{\tau_{0}}^{\tau}=\frac{1}{2(\epsilon-1)}\ln{\bigg{[}\frac{1}{2}\bigg{(}1+\bigg{(}\frac{\tau}{\tau_{0}}\bigg{)}^{2}\bigg{)}\bigg{]}},$ (83)
    for $N_{c}<N<N_{b}$ where the bouncing phase ends at $N=N_{b}$, and $N_{0}$ and $\tau_{0}$ are reference variables for the e-folding and conformal time such that $a(\tau_{0})=a_{0}$ in Eqn. (3). Using this we can express the conformal Hubble parameter in terms of the e-folds as follows:
    ${\cal H}(N)$ $=$ $\frac{1}{\tau_{0}(\epsilon-1)}\frac{\tau}{\tau_{0}}\bigg{(}1+\bigg{(}\frac{\tau}{\tau_{0}}\bigg{)}^{2}\bigg{)}^{-1},$ (84)
    $=$ $\frac{1}{2\tau_{0}(\epsilon-1)}\exp{[-2(N-N_{0})(\epsilon-1)]}\sqrt{2\exp{[2(N-N_{0})(\epsilon-1)]}-1}.$
    Continuing from here we can get the relation for $\epsilon(N)$ using its definition, and the relation in Eqn ((b)), as follows:
    $\epsilon(N)=1-\frac{{\cal H}^{\prime}}{{\cal H}^{2}}=1-\frac{(\epsilon-1)(\tau_{0}^{2}-\tau^{2})}{\tau^{2}}=\epsilon+\frac{(1-\epsilon)}{2\exp{[2(N-N_{0})(\epsilon-1)]}-1},$ (85)
    which tells us that for large enough $N$ the exponential in the denominator becomes large enough so that $\epsilon$ again becomes constant in the respective phase. The $\epsilon$ values on the RHS take again similar values during the ekpyrotic $(\epsilon>3)$ or matter $(\epsilon=3/2)$ bouncing phase. Finally, from the definition of the parameter $\eta(N)$ and Eqn. ((b)), we have the following relation:
    $\eta(N)=\frac{\epsilon^{\prime}}{\epsilon{\cal H}}$ $=$ $2(\epsilon-1)^{2}\times\frac{\tau_{0}^{2}}{\tau^{2}}\frac{1+\big{(}\frac{\tau}{\tau_{0}}\big{)}^{2}}{1+\epsilon\big{(}\frac{\tau^{2}}{\tau_{0}^{2}}-1\big{)}},$ (86)
    $=$ $2(\epsilon-1)^{2}\times\frac{1}{(2\exp{[2(N-N_{0})(\epsilon-1)]}-1)}\frac{2\exp{[2(N-N_{0})(\epsilon-1)]}}{1+2\epsilon(\exp{[2(N-N_{0})(\epsilon-1)]}-1)}.$
  3. (c)
    Phase-III  (First Slow-Roll): We can now begin analysing the slow-roll parameters in the phases during inflation starting with the first slow-roll. We assume that throughout the phase the SR parameter $\eta(N)$ remains a small and negative almost constant value. This further leads us to conclude from its definition that the first SR parameter $\epsilon(N)$ is also a constant since:
    $\eta$ $=$ $\frac{\epsilon^{\prime}}{\epsilon{\cal H}}=\frac{1}{\epsilon}\frac{d\epsilon}{dN},$ (87)
    and that it also remains an almost constant value during $N_{*}<N<N_{s}$ to satisfy the assumption of $\eta(N)$. The value $N_{*}$ here refers to a reference e-folding value corresponding to the start of SRI, which we choose to be $\epsilon(N=N_{*})={\cal O}(10^{-3})$. $N_{s}$ refers to the point where SRI ends and the next USR phase begins.
  4. (d)
    Phase-IV  (Ultra Slow-Roll): Here we assume that the transition from the SRI to the USR occurs sharply such that the value of the second SR parameter suddenly changes to become, $\eta(N)\sim{\cal O}(-6)$, which is a negatively large value. It stays the same throughout the interval of this phase, $N_{s}<N<N_{e}$, and, using its definition, we can write the following relation for the first SR parameter:
[…]
    which tells us that $\epsilon(N)$ suffers an exponential decrease during the USR phase. The value $\epsilon(N_{s})$ matches its value during the end of the SRI phase.
  5. (e)
    Phase-V  (Second Slow-Roll): The previous assumption regarding the nature of the transition between the SRI and USR phases follows again after we exit the USR and enter the final SRII phase. The second SR parameter, this time, increases from $\eta(N)\sim{\cal O}(-6)$ to $\eta(N)\sim{\cal O}(-1)$, indicating the end of inflation. This value remains the same throughout this phase until the first SR parameter climbs back close to unity. Its behaviour can be understood in a similar fashion done before as follows:
[…]
    where we take $\epsilon(N_{\rm end})\sim{\cal O}(1)$ as the moment when inflation truly ends with final e-foldings being $N_{\rm end}$. As usual, continuity demands that the parameters remain the same value at the transition instant $N=N_{e}$.
Appendix B Comoving curvature perturbation mode solutions